OpenStudy (anonymous):

solve the inequality. explain all steps 3a−4≤5

OpenStudy (texaschic101):

step by step... 3a - 4 <= 5 (add 4 to both sides) 3a - 4 + 4 <= 5 + 4 (combine like terms) 3a <= 9 (divide both sides by 3) 3a/3 <= 9/3 a <= 3 any questions ?